# Explainability of AI Uncertainty: Application to Multiple Sclerosis Lesion Segmentation on MRI

### v1

@@ 4.3 Feature importance analysis
Figure 5 presents coefficients of the Elastic-Net model fitted to the entire feature space, ranked by magnitude; 20 coefficients explaining the biggest part of variance are displayed. Since all the features were standardized before model fitting, the absolute values are comparable and indicate feature importance. The features with the highest absolute importance explain the most variability in uncertainty. Positive values coefficients indicate features that correlate with increased uncertainty (when the feature value rises, uncertainty typically increases), while negative coefficients represent features associated with decreased uncertainty.
Figure 5: Barplot with linear explainer coefficients treated as feature importance. Barplots are built for Train, Test-in, and Test-out sets. Positive values - direct relationship with uncertainty, and negative values - direct relationship with certainty. Twenty features with the highest absolute coefficients are built.
@@ 4.4 Selected features analysis
Figure 7: Lesion examples sampled based on the feature distribution. The shape, location, and intensity features were selected based on the feature importance analysis. CLs were obtained based on 0, 25, 50, 75, and 100 percentiles (columns from left to right) of each feature distribution on a particular set (rows - Train, Test-in, or Test-out). All the features have a positive relationship with lesion-wise uncertainty, except for GM overlap and lesion intensity mean absolute deviation, which is negatively related to uncertainty. Only lesions with a non-zero overlap with the ground truth are displayed.
@@ 4.5 Expert rater questionnaire
Figure 8 summarizes the results of the expert rater questionnaire, identifying key factors contributing to the low visibility and reduced manual segmentation quality of CLs on MP2RAGE. The proposed concepts, were additioned by the respondents with respect to the ”Location in the brain” category. The proposed concepts include i) similarity in appearance to perivascular spaces in the insular cortex and ii) temporal lobes.
The responses indicated that subpial lesions, small lesions, lesions with low contrast with the surrounding tissue, and scans with low resolution are the most challenging to annotate (4 out of 5 raters). Lesion location relative to the cortex was also selected by the majority of participants, suggesting that both lesions with unclear cortical involvement and intracortical lesions pose significant challenges for annotators. Image quality-related factors such as artifacts and blurriness were also selected by the majority of respondents as additional obstacles. Furthermore, the presence of high-intensity values or heterogeneous intensity patterns within a lesion was noted. Two experts specifically highlighted the insular cortex and temporal lobes as brain regions where CLs are more difficult to annotate. None of the respondents selected patient-related factors as influencing annotation difficulty, and all reported that patient-specific information was not accessed before annotation.
@@ 5.1 Information carried by instance-wise uncertainty
This alignment between model uncertainty and annotator uncertainty suggests that deep ensemble uncertainty also encodes aleatoric noise. Deep ensemble uncertainty is commonly associated with epistemic uncertainty and, thus, more strongly correlated with model error than other UQ methods [37, 70]. However, this assumptions stems largely from low-noise tasks. In contrast, CL segmentation is inherently noisy. Moreover, hidden data biases, such as underrepresentation of certain lesion subtypes (e.g., subpial or small intracortical lesions) [30, 71], may also contribute to increased uncertainty. This interplay of factors raises questions about the separability of the sources of aleatoric and epistemic uncertainty [72].
As discussed in Section 4.4, non-linear dependencies are present between uncertainty and selected features with high importance. Partially, an unexplained uncertainty should be attributed to the non-linear relationship. However, modelling these dependencies requires more data, since random forests explainers did not outperform linear models in goodness of fit, likely due to overfitting in small datasets [73]. While linear models are pragmatic for limited data, their results require supplementary analysis (e.g., partial dependence plots) to characterize non-linear dependencies.
@@ 5.2 Explainability of uncertainty under domain shifts
We evaluated uncertainty and the proposed framework performance under domain shift — a clinically realistic scenario where prediction quality varies across settings. As shown in Section 4.2, the explanatory power of uncertainty degraded progressively from Train to Test-in to Test-out, indicating reduced informativeness with growing domain shift. We hypothesize that this degradation is caused by changes in relevant features and a deterioration in uncertainty quality. Prior work already showed that UQ reliability declines under distributional shifts [70, 54, 48].
@@ 5.3 Predictive uncertainty and annotator confidence
However, a direct one-to-one mapping between clinically identified factors and engineered lesion features is not feasible. The importance of lesion location, as emphasized by experienced raters, also extends to predictive uncertainty. The non-linear relationship between lesion uncertainty and GM overlap could be explained by the fact that the model is both uncertain on lesions with unclear cortex involvement (GM overlap 0) and on the intracortical lesions (GM overlap = 1). Translating engineering characteristics into understandable concepts is not straightforward [74] and beyond the scope of this work.
@@ 5.4 Future implications of uncertainty explanations
Explainability of uncertainty has practical values in many different domains. From an engineering perspective, characterizing difficult and ambiguous cases may contribute to the development of methods to improve model robustness or uncertainty calibration. From the stakeholder perspective, it supports human-centered AI development by enabling transparency, fostering trust, and integrating expert feedback.
On the one hand, explaining uncertainty and visualizing important features fosters communication with medical professionals. On the other hand, expert feedback enriches our understanding of model behavior, revealing its data-dependent nature.
Our results suggests that unexplained uncertainty is partially attributed to the approximation error. This means that some uncertainty values are misleading and have the power to bias medical decision-making into the wrong direction. Quantifying this unexplained component could support trust calibration, aligning human trust with actual model reliability. However, further work is needed to evaluate how clinicians perceive and utilize such explanations.
While our framework was demonstrated in the context of MRI-based lesion segmentation, it is easily adaptable to other UQ methods, machine learning tasks (e.g., classification), imaging modalities, and clinical applications.
Finally, while the proposed analysis is correlation-based, it makes hypotheses about the causal relationship, which should be further investigated in future work.
@@ 6 Conclusion
We proposed a novel framework for explaining AI-models predictive uncertainty, beyond classical error estimation analysis. Our results on lesion uncertainty confirm the relationship between uncertainty and error and link high uncertainty with clinically relevant concepts. We also show the limitation of UQ in the out-of-domain setting and assess the presence of misleading uncertainty values. Overall, increasing the understanding of AI model reliability in medical image analysis is needed to increase transparency, facilitate the human-AI interactions, and ultimately support building stakeholders trust in AI decisions.
@@ Acknowledgment
This work was supported by the Hasler Foundation Responsible AI program (MSxplain) and the Research Commission of the Faculty of Biology and Medicine (CRFBM) of UNIL. This research was partially supported by the Intramural Program of NINDS, NIH, USA. We acknowledge access to the facilities and expertise of the CIBM Center for Biomedical Imaging, a Swiss research center of excellence founded and supported by Lausanne University Hospital (CHUV), University of Lausanne (UNIL), École polytechnique fédérale de Lausanne (EPFL), University of Geneva (UNIGE) and Geneva University Hospitals (HUG). The authors wish to thank Dr. Mara Graziani for valuable consultations and discussions at the initial stage of this research.
@@ Appendix A Metrics for model performance evaluation
In the case of segmentation, the F1 score is called the Dice similarity score; in case of lesion detection, the F1, precision, and recall are called, LF1, LPPV, and LTPR, respectively.
@@
$F_{1}=\frac{2TP}{2TP+\kappa\cdot FP+FN},\quad\kappa=h(r^{-1}-1)$
@@ Appendix B Adjusted intersection over the union definition
During the preparation of this work the authors used Grammarly and GhatGPT for grammatical and stylistical errors correction. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.

### v2

@@ 4.3 Feature importance analysis
Figure 5 presents coefficients of the Elastic-Net model fitted to the entire feature space, ranked by magnitude; 20 coefficients explaining the biggest part of the variance are displayed. Since all the features were standardized before model fitting, the absolute values are comparable and indicate feature importance. The features with the highest absolute importance explain the most variability in uncertainty. Positive values of coefficients indicate features that correlate with increased uncertainty (when the feature value rises, uncertainty typically increases), while negative coefficients represent features associated with decreased uncertainty.
Figure 5: Barplot with linear explainer coefficients treated as feature importance. Barplots are built for Train, Test-in, and Test-out sets. Positive values - direct relationship with uncertainty, and negative values - direct relationship with certainty. Twenty features with the highest absolute coefficients are built. MAD -mean absolute deviation.
@@ 4.4 Selected features analysis
Figure 7: Lesion examples sampled based on the feature distribution. The shape, location, and intensity features were selected based on the feature importance analysis. CLs were obtained based on 0, 25, 50, 75, and 100 percentiles (columns from left to right) of each feature distribution on a particular set (rows - Train, Test-in, or Test-out). All the features have a positive relationship with lesion-wise uncertainty, except for GM overlap and lesion intensity mean absolute deviation, which are negatively related to uncertainty. Only lesions with a non-zero overlap with the ground truth are displayed.
@@ 4.5 Expert rater questionnaire
Figure 8 summarizes the results of the expert rater questionnaire, identifying key factors contributing to the low visibility and reduced manual segmentation quality of CLs on MP2RAGE. The proposed concepts were additioned by the respondents with respect to the ”Location in the brain” category. The proposed concepts include i) similarity in appearance to perivascular spaces in the insular cortex and ii) temporal lobes.
The responses indicated that subpial lesions, small lesions, lesions with low contrast with the surrounding tissue, and scans with low resolution are the most challenging to annotate (4 out of 5 raters). Lesion location relative to the cortex was also selected by the majority of participants, suggesting that both lesions with unclear cortical involvement and intracortical lesions pose significant challenges for annotators. Image quality-related factors, such as artifacts and blurriness, were also selected by the majority of respondents as additional obstacles. Furthermore, the presence of high-intensity values or heterogeneous intensity patterns within a lesion was noted. Two experts specifically highlighted the insular cortex and temporal lobes as brain regions where CLs are more difficult to annotate. None of the respondents selected patient-related factors as influencing annotation difficulty, and all reported that patient-specific information was not accessed before annotation.
@@ 5.1 Information carried by instance-wise uncertainty
This alignment between model uncertainty and annotator uncertainty suggests that deep ensemble uncertainty also encodes aleatoric noise. Deep ensemble uncertainty is commonly associated with epistemic uncertainty and, thus, more strongly correlated with model error than other UQ methods  (Lakshminarayanan et al., 2017; Ovadia et al., 2019). However, this assumption stems largely from low-noise tasks. In contrast, CL segmentation is inherently noisy. Moreover, hidden data biases, such as underrepresentation of certain lesion subtypes (e.g., subpial or small intracortical lesions)  (Cagol et al., 2024; Beck et al., 2022), may also contribute to increased uncertainty. This interplay of factors raises questions about the separability of the sources of aleatoric and epistemic uncertainty  (Valdenegro-Toro and Saromo, 2022).
As discussed in Section 4.4, non-linear dependencies are present between uncertainty and selected features with high importance. Partially, an unexplained uncertainty should be attributed to the non-linear relationship. However, modelling these dependencies requires more data, since random forest explainers did not outperform linear models in goodness of fit, likely due to overfitting in small datasets  (Hastie et al., 2009). While linear models are pragmatic for limited data, their results require supplementary analysis (e.g., partial dependence plots) to characterize non-linear dependencies.
@@ 5.2 Explainability of uncertainty under domain shifts
We evaluated uncertainty and the proposed framework performance under domain shift — a clinically realistic scenario where prediction quality varies across settings. As shown in Section 4.2, the explanatory power of uncertainty degraded progressively from Train to Test-in to Test-out, indicating reduced informativeness with growing domain shift. We hypothesize that this degradation is caused by changes in relevant features and a deterioration in uncertainty quality. Prior work already showed that UQ reliability declines under distributional shifts  (Ovadia et al., 2019; Malinin et al., 022a, 022b).
@@ 5.3 Predictive uncertainty and annotator confidence
However, a direct one-to-one mapping between clinically identified factors and engineered lesion features is not feasible. The importance of lesion location, as emphasized by experienced raters, also extends to predictive uncertainty. The non-linear relationship between lesion uncertainty and GM overlap could be explained by the fact that the model is both uncertain on lesions with unclear cortex involvement (GM overlap 0) and on the intracortical lesions (GM overlap = 1). Translating engineering characteristics into understandable concepts is not straightforward  (Gerdes, 2024) and beyond the scope of this work.
@@ 5.4 Future implications of uncertainty explanations
Understanding and explaining uncertainty holds practical importance across a range of applications. From a technical standpoint, identifying ambiguous or difficult cases can guide efforts to improve model reliability and calibration (Lambert et al., 2024). For clinical users and other stakeholders, such explanations aim to enhance the transparency and support more informed interactions with automated decision-support tools (Lekadir et al., 2025). By relating uncertainty to clinically meaningful features and presenting this information through visualizations, the proposed framework facilitates the interpretation of model behavior by medical professionals. In parallel, structured expert feedback enhances our understanding of how models perform under different data conditions, identifying areas where model outputs may be less reliable or misunderstood.
Our findings indicate that a portion of predictive uncertainty cannot be explained by the examined features and may arise from UQ methods limitations and approximation errors. Such unexplained uncertainty is potentially uninformative and, thus, has the potential to mislead annotator decisions or impede their confidence. Further efforts in identifying the failures in UQ will contribute to a better trust calibration, aligning perceived and actual reliability (Zhang et al., 2020).
Although this study is centered on lesion segmentation in brain MRI, the framework is broadly applicable to other types of uncertainty, learning tasks (e.g., classification), imaging modalities, and clinical scenarios. Finally, while the current analysis is correlational, future work should aim to explore underlying causal relationships to improve the explanatory power and practical utility of uncertainty estimates.
@@ 6 Conclusion
This study addresses a core challenge in the development of trustworthy AI for medical imaging: how to make uncertainty estimates understandable and clinically meaningful. Motivated by the need for transparency, reliability, and trust in high-risk settings such as healthcare, we proposed a framework that links lesion-level uncertainty to patient- and lesion-specific characteristics, moving beyond traditional error-based evaluation.
Our results show that uncertainty carries clinically relevant information and reveals its sensitivity to domain shift. We also identify cases of unexplained uncertainty that could mislead users, underscoring the importance of understanding not just the presence of uncertainty, but its implications. By facilitating clearer communication and supporting clinician feedback, this framework contributes to building trust in automated systems. It offers a practical step toward integrating uncertainty-aware AI tools in real-world clinical settings and provides a foundation for future studies across different modalities and tasks.
@@ Acknowledgment
This work was supported by the Hasler Foundation Responsible AI program (MSxplain) and the Research Commission of the Faculty of Biology and Medicine (CRFBM) of UNIL. This research was partially supported by the Intramural Program of NINDS, NIH, USA. We acknowledge access to the facilities and expertise of the CIBM Center for Biomedical Imaging, a Swiss research center of excellence founded and supported by Lausanne University Hospital (CHUV), University of Lausanne (UNIL), École polytechnique fédérale de Lausanne (EPFL), University of Geneva (UNIGE), and Geneva University Hospitals (HUG). The authors wish to thank Dr. Mara Graziani for valuable consultations and discussions at the initial stage of this research.
@@ Appendix A Metrics for model performance evaluation
In the case of segmentation, the F1 score is called the Dice similarity score; in the case of lesion detection, the F1, precision, and recall are called LF1, LPPV, and LTPR, respectively.
@@
$nDSC=\frac{2TP}{2TP+\kappa\cdot FP+FN},\quad\kappa=h(r^{-1}-1)$
@@ Appendix B Adjusted intersection over the union definition
During the preparation of this work, the authors used Grammarly and ChatGPT to correct grammatical and stylistic errors. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.